OpenStudy (dedi):

A bungee jumper falls 97.2 feet before bouncing back up at the end of his bungee cord. Then he falls 64.8 feet before bouncing up again. On his third descent, he falls 43.2 feet before going back up. What is the total distance the bungee jumper drops after five falls if the distances continue in this pattern? (Do not include the distance traveled up.) A)234.0 ft B)253.2 ft C)789.8 ft D)1,281.8 ft

OpenStudy (amistre64):

hmm, what models the pattern best? what type of sequence?

OpenStudy (dedi):

Geometric sequence

OpenStudy (amistre64):

lets chk that if its geometric, then we have a common ratio 97.2/64.8 = 64.8/43.2 which is good, so we know our r value

OpenStudy (amistre64):

given a starting value a1, and a common ratio , r we can use the formula:\[S_n=a_1\frac{1-r^n}{1-r}\]

OpenStudy (dedi):

okay is it d?

OpenStudy (amistre64):

not when we use the formula, i dont get d

OpenStudy (amistre64):

define the varaibles for me: a1, r and n

OpenStudy (dedi):

I think its 97.2,64.8

OpenStudy (amistre64):

a1 is 97.2 and since a1 r = a2; r = a2/a1 and n is stated for 5

OpenStudy (dedi):

is it b or c?

OpenStudy (amistre64):

lets use the formula ....

OpenStudy (amistre64):

what is that value of r, since we have already defined a1 and n